Find the percent by mass of oxygen (O) in potassium chlorate, KClO3.


1
Expert's answer
2022-03-04T13:20:02-0500

As we know that , the molar mass of KClO3=39+35.5+48g =122.5g

So, the percentage by mass of oxygen(O)

="\frac{48}{122.5}×100" = 39.18%